-16n+12=39-7n
My question is what is the rule for combining the variable.
If I +7 I get -11 but if I =16 I then get +11
There must be a rule I don't know...please help!!

Answer by Timnewman(323)

Hi dear!
The rule of bringing these variables states that if the the variable carries positive sign (+),subtract (-) the variable from both side;and if the variable carries nagative sign,add the variable to both side of the equation.
Applying the rule to the solution of your problem,
-16n+12=39-7n
since -7n is nagative,add 7n to both sides
7n-16n+12=39-7n+7n
Your equation becomes
-8n+12=39
also,since 12 is positive,subtract it from bothsides
-8n+12-12=39-12
Which becomes
-8n=27
IS THAT OK?